### 07 Apr 2026 v2

# Temporal Variation of the Coronal Radius Parameter in a Jetted Tidal Disruption Event: Swift J1644+57

Abstract
Tidal Disruption Events are exotic astrophysical phenomena where matter from a star or the interstellar medium is captured by a supermassive black hole. The process liberates enormous energy, within a few months to a year timescale, enough to detect dormant black holes in near as well as the farthest galaxies. We revisit the long-term spectral variabilities associated with the jetted Tidal Disruption Event Swift J1644+57 by exploring the archival X-ray data obtained with Swift/XRT and XMM-Newton observatories. Our analysis reveals that the spectral indices decrease non-monotonically as Swift J1644+57 evolves with time. We also find that the soft (0.3-1.5 keV) and hard (1.5-10 keV) X-ray photon counts are highly correlated with a maximum correlation coefficient of 0.95 and peak at zero lag. Moreover, the soft and hard band variabilities obtained from XMM-Newton observations are highly correlated with a Pearson cross-correlation coefficient of 0.96. This indicates that the soft and hard X-ray photons are emitted from the same site, which is most likely a Compton cloud, i.e., the corona. Assuming the hard X-ray photons originate from the corona, we find that the coronal parameter undergoes rapid expansion during the early phases when accompanied by a relativistic jet launching and subsequently evolves toward a state of saturation with minor fluctuations in the latter stages. The temporal variation of the coronal radius parameter ($R_{cor}$) is consistent with a simple theoretical conjecture. We also discuss the application of our analytical outcomes to other jetted and non-jetted tidal disruption events.
Accretion — Black Hole Physics — Tidal Disruption Events — Radiative processes–Individual: Swift J1644+57
I Introduction
Most galaxies are thought to harbor supermassive black holes (SMBHs) within the mass range of $10^{6}\,M_{\odot}\lesssim{M_{\rm bh}}\lesssim 10^{10}\,M_{\odot}$ (Kormendy and Richstone, 1995; Kormendy and Ho, 2013). The mass of the SMBH has been estimated using the reverberation mapping technique (Peterson et al., 2004), measuring the proper motion of stars orbiting around the SMBH, and the $M-\sigma_{*}$ relation, where the velocity dispersion of stars ($\sigma_{*}$) around the galactic bulge is correlated with the mass of the SMBH ($M$).
Tidal disruption events (TDEs) can provide a unique way to probe the dormant SMBHs in inactive galaxies. A star is tidally disrupted when it reaches within the tidal disruption radius $R_{\rm t}=(M_{\rm bh}/m_{*})r_{*}$, where $M_{\rm bh}$ is the mass of the SMBH, $m_{*}$ is the stellar mass, and $r_{*}$ is the stellar radius.
The tidal disruption radius is expressed by
$\frac{R_{\rm t}}{r_{\rm g}}\approx 50\,\left(\frac{M_{\rm bh}}{10^{6}\,M_{\odot}}\right)^{-2/3}\left(\frac{m_{*}}{\,M_{\odot}}\right)^{-1/3}\left(\frac{r_{*}}{\,R_{\odot}}\right)$ (1)
in units of the gravitational radius, $r_{\rm g}=GM_{\rm bh}/c^{2}$, where is $G$ is the gravitational constant, $M_{\rm bh}$ is the mass of the black hole, and $c$ is the speed of light.
After tidal disruption of the star, the stellar debris falls toward the SMBH and releases its gravitational energy as radiation, generating a luminosity large enough to be comparable to the Eddington limit. According to the standard TDE theory (Rees 1988; see also Rossi et al. 2021 for a review), the resultant luminosity decays with time, following the power law of time $t^{-5/3}$, for a span of a few 100 days over the wide range of optical/UV to X-ray wavebands. It is a key observational feature that the light curve decays with $t^{-5/3}$. Though a significant number of TDEs can deviate from the standard $t^{-5/3}$ decay rate due to stellar internal structure (Lodato and Rossi, 2011), orbital eccentricity (Hayasaki et al., 2013, 2016, 2018; Park and Hayasaki, 2020; Zhong et al., 2023), and/or the penetration factor which is the ratio of the tidal disruption to pericenter radii (Guillochon and Ramirez-Ruiz, 2013).
The first observational evidence of TDEs dates back to the 1990s (Bade et al., 1996) by the All-Sky Survey of the ROSAT satellite. Since then, through multi-wavelength campaigns, $\sim 100$ (Gezari, 2021) have been discovered in a wide range of wavebands from radio to optical/UV to X-ray bands. While most TDEs have been observed in optical/UV wavelengths, $\sim 10$ TDEs have been detected simultaneously in the UV/optical and soft X-ray wavelengths. In addition to these thermal TDEs without a relativistic jet (so-called non-jetted TDEs), there have been so far four remarkable TDEs that accompany a relativistic jet (jetted TDEs), such as Swift J164449.3+573451 (hereafter, Swift J1644+57) (Levan et al., 2011; Bloom et al., 2011a, b; Burrows et al., 2011), Swift J2058+05 (Cenko et al., 2012), Swift J1112-8238 (Brown et al., 2015), and AT2022cmc (Andreoni et al., 2022).
The event rate of non-jetted TDEs is estimated to be $1\times 10^{-5}\,{\rm yr^{-1}}$ per galaxy for a typical galaxy (Donley et al., 2002; Magorrian and Tremaine, 1999; Wang and Merritt, 2004), whereas the event rate of jetted TDEs is estimated $0.03~{\mathrm{Gpc}}^{-3}{\mathrm{yr}}^{-1}$ (Sun et al., 2015; De Colle and Lu, 2020; Andreoni et al., 2022). Recent observations indicate that many Changing look active galactic nuclei (AGNs) have been associated with TDEs (Eracleous et al., 1995; Merloni et al., 2015; Ricci et al., 2020, 2021; Jana et al., 2021). This could also potentially increase the population of TDEs with e-ROSITA (Merloni et al., 2012) in the future. These event rates indicate the remarkable scarcity of the jetted TDE population. However, the population of jetted TDEs estimated above indicates a relatively small beaming angle ($\leq$ 1°). Of late, the discovery of off-axis jetted TDEs, such as IGR J12580+0134 in the nucleus of NGC 4845 galaxy (Walter et al., 2011) and the subsequent radio emission (Lei et al., 2016) suggest that the TDEs, which emit hard X-rays with weak radio emission, can show a global property of X-ray TDEs. Several non-jetted, radio-emitting TDEs have also been detected (Alexander et al., 2020; Cendes et al., 2021). Moreover, the recently discovered association of astrophysical IceCube neutrinos with three TDE candidates (Stein et al., 2021; Hayasaki, 2021) invoked a new era of multi-messenger astronomy of TDEs.
There are differences in X-ray spectral hardness between jetted TDEs and non-jetted TDEs. Non-jetted TDEs often exhibit softer spectra, where photon index $\Gamma>2.5$ is associated with a thermal component (Komossa, 2017). Contrary to that, jetted TDEs have much harder spectra with $\Gamma\leq 2$. In addition, no clear evidence of a thermal component could be observed from the X-ray spectral energy distribution (Mangano et al., 2016; Seifina et al., 2017). The domain of jetted TDEs began in March 2011 with the discovery of Swift J1644+57. The origin of high-energy $\gamma$-rays is traced back to a galaxy at $z=0.35$ (Levan et al., 2011). The peak X-ray luminosity exceeded the Eddington limit with the inferred black hole mass $3\times 10^{6}M_{\odot}$ (Bloom et al., 2011b; Burrows et al., 2011; Levan et al., 2011). The source was detected in radio (Cendes et al., 2014; Berger et al., 2012; Zauderer et al., 2013), infra-red (Wiersema et al. 2012); X-rays (Reis et al., 2012; Saxton et al., 2012; González-Rodríguez et al., 2014; Mangano et al., 2016; Zauderer et al., 2013), $\gamma$-rays (Bloom et al., 2011b; Burrows et al., 2011). Unlike X-ray luminosity, which decayed over the next 500 days, the radio luminosity increased (Berger et al., 2012; Zauderer et al., 2013) as the event progressed.
Figure 1: Upper Panel: X-ray light curves of Swift J1644+57, obtained from the Swift/XRT, are presented for 0.3–1.5 keV (square-light-salmon) and 1.5–10 keV energy bands (diamond-greenish-blu-e points). Lower panel: $\zeta$-DCF cross-correlation between 0.3-1.5 keV and 1.5–10 keV ranges are presented for Swift J1644+57 using Swift/XRT data. Both the energy bands are highly correlated ($\rho_{max}=0.95$). Considering the error bars, no delay is observed between the two X-ray bands.
Seifina et al. (2017) showed that the spectral hardness of the source evolved from $\Gamma\sim 1.8$ to $\Gamma\sim 1.2$ for the first 200 days since the outburst. However, the variation of spectral index met sudden fluctuations associated with the dips found in the light curve (Mangano et al., 2016). From the X-ray spectral shape, luminosity, and the absence of the thermal component, Zauderer et al. (2013) asserted that the X-ray photons of Swift J1644+57 could not have originated from the accretion disk. Rather, the origin of the X-rays could be from the internal dissipation within the jet at a large distance ($r\sim 10^{4}-10^{5}r_{\rm g}$) from the SMBH. Contrary to their idea, Lei et al. (2016) suggested the X-ray emission from the off-axis jetted-TDE candidate IGR J12580+0134 originates from the disk-corona. Beamed accretion, much like in the case of Blazers, could stand out as a possible scenario for the hyper accretion rate. However, in the case of Swift J1644+57, the luminosity kept falling with decreasing X-ray photon index ($\Gamma$). This observed correlation between luminosity and $\Gamma$ opposes the “harder when brighter” (Krawczynski et al., 2004; Gliozzi et al., 2006; Zhang et al., 2006; Pandey et al., 2017) nature of the X-ray bright Blazars and aligns with the accretion disk dominated Seyferts or galactic black holes (GBHs) (e.g., Yang et al., 2015; Jana, 2022; Jana et al., 2026). Stronger evidence of the accretion disk origin of the X-ray emission in Swift J1644+57 came from Kara et al. (2016), where reverberation lag was observed.
A quasi-periodic oscillation (QPO) of 200 s, within the first 20 days from the peak luminosity, was discovered from the X-ray light curves in the 2-10 keV energy band, using the Suzaku and XMM-Newton observations (Reis et al., 2012; Saxton et al., 2012). Similar QPOs in X-rays have been observed for two other TDE candidates, namely 2XMM J123103.2+110648 (Lin et al., 2013) and ASASSN-14li (Pasham et al., 2019). Nikołajuk and Walter (2013) suggested the detection of a QPO within 0.00080.0008-$0.004$  Hz for IGR J12580+0134, where the timescale is consistent for a TDE with central black hole mass $\sim 10^{5}M_{\odot}$. There are many arguments that the QPOs can bridge the mass gap between the accretion processes of galactic, stellar-mass black holes, and SMBHs (Vaughan and Uttley, 2005). However, very few of them are detected with statistical confidence for AGNs (Gierliński et al., 2008; Alston et al., 2015; Ashton and Middleton, 2021). There are many challenges for observing QPOs originating from AGNs, apart from achieving a statistically significant signal to noise ratio (Vaughan and Uttley, 2006). Considering the mass of Swift J1644+57, the 200 s QPO found in the X-ray light curve requires a scale of $\mathcal{O}(10)\,r_{\rm s}$ for a non-rotating SMBH. This scale fits within the range of the typical coronal radius parameter (hereafter $R_{cor}$) of TDEs as proposed by Mummery and Balbus (2021), and is also consistent with the corona size, which is observationally suggested for many AGNs (Alston et al., 2020). Thus, we expect to find a reasonable explanation of hard X-ray variations through the disk-corona geometry by analyzing the existing data of jetted TDEs. The most promising mechanism for the luminous emission of TDEs is a thermal blackbody emission. Gezari (2021) suggested that there is a bimodal distribution of the TDE thermal emission, indicating that the soft X-ray photons are produced from the inner disk region, whereas the optical/UV photons could be emitted from the photospheric outflows. The thermal photons produced from the disk can be a good candidate source for the Compton scattering. This suggests a disk-corona system can exist in X-ray TDEs.
[…]
In this paper, we study how the Compton cloud of the jetted TDE Swift J1644+57 varies with time by using the Optxagnf model and whether the model can explain the hard X-ray properties. The paper is structured in the following way. In Section 2, we explain the data analysis method and the subsequent processing. In Section 3, we describe our results obtained from data analysis. We then discuss our findings and the implications on TDEs in Section 4. Finally, we summarize our conclusions in Section 5.
II Data Analysis Process
We used the archival data of XMM-Newton and Swift/XRT obtained through HEASARC [^1] http://heasarc.gsfc.nasa.gov/. We reprocessed all data using HEAsoft v6.26.1 (Arnaud, 1996), which includes XSPEC v12.10.1f. We use the following cosmological parameters in this work: $H_{0}$ = 70 km s-1 Mpc-1, $\Lambda_{0}$ = 0.73, $\Omega_{M}$ = 0.27 (Bennett et al., 2003). With the cosmological parameters, the luminosity distance of Swift J1644+57 was considered 1402.8 Mpc (Levan et al., 2011).
II.1 Swift/XRT
The X-ray telescope (XRT) (Burrows et al., 2005) on board Swift has the highest cadence monitoring observation of Swift J1644+57 in the X-ray band (0.3 to 10.0 keV), especially during the TDE from 2011. Depending on the brightness of the source, XRT observed this source in both Window Timing (WT) and Photon Counting (PC) modes. The source was observed by Swift/XRT over $\sim$500 times from 2011-03-31 to 2012-09-30. We stacked all observations into 17 segments depending on the brightness of the source. When the event started, the source was brighter. Thus, we stacked $\sim 15$ observations into a single spectrum and produced the combined spectrum. This procedure is followed for the first $154$ observations. We stacked these observations into 10 combined observations (XRT1 to XRT10) (details are given in Table 1). After that, the source became dimmed with respect to previous observations. For that, we stacked $\sim 40$ observations into a single spectrum. This process is followed for the next 250 observations and produces the stacked spectrum from XRT11 to XRT15 (see Table 1). After 2012-03-17, we stacked 176 observations and 108 observations into a combined observation for XRT16 and XRT17, respectively. We use the online tool “XRT product builder” [^2] http://swift.ac.uk/user_objects/ (Evans et al., 2009) to extract the spectrum and light curves. This product builder performs all necessary processing and calibration and produces the final spectra and light curves of Swift J1644+57  WT and PC modes. The GRPPHA task is used with 10 counts per bin for XRT spectra. For more details, see Table 1.
II.2 XMM-Newton
We reprocessed the XMM-Newton Epic-pn data using the Science Analysis Software (SAS) v17.0.0 and updated calibration files. We reprocessed the data with the “epproc” task. For each observation ID, we created good time interval (GTI) files after eliminating the intervals of flaring particle background based on the count rate in the light curves above 10 keV. We filtered the event lists to get the good events with a pattern <=<=4 for a given GTI file and generated the cleaned event lists. We extracted source and background spectra using circular regions of radii 30 and 50 arcsec, respectively. We generated the light curves in 0.220.2-2 keV and 3103-10 keV ranges with 10 s time bin and corrected from effects like vignetting, bad pixels, and variations in PSF. We also checked for any pile-up effects in the data. We generated the redistribution matrix and ancillary response files using the SAS tasks “rmfgen” and “afrgen”, respectively. We grouped the spectrum with a minimum of 20 counts per bin and oversampled it by a factor of 3 using the “specgroup” task.
Table 1: Observation log for Swift/XRT and XMM-Newton observations of Swift J1644+57.
ID Date Obs. ID Instrument Total Exposure
(yyyy-$m$-$d$) (ks)
XRT1 2011-03-31 00031955002 Swift/XRT 224
-2011-04-14 -00031955016
XMM1 2011-03-31 0658400701 XMM-Newton/Epic-pn 26
XRT2 2011-04-15 00031955017 Swift/XRT 210
-2011-04-30 -00031955032
XMM2 2011-04-16 0678380101 XMM-Newton/Epic-pn 25
XMM3 2011-04-30 0678380201 XMM-Newton/Epic-pn 29
XRT3 2011-05-01 00031955033 Swift/XRT 137
-2011-05-13 -00031955047
XRT4 2011-05-14 00031955048 Swift/XRT 162
-2011-05-29 -00031955063
XMM4 2011-05-16 0678380301 XMM-Newton/Epic-pn 30
XRT5 2011-05-30 00031955064 Swift/XRT 125
-2011-06-12 -00031955078
XMM5 2011-05-30 0678380401 XMM-Newton/Epic-pn 30
XRT6 2011-06-13 00031955079 Swift/XRT 131
-2011-06-26 -00031955093
XRT7 2011-06-27 00031955094 Swift/XRT 142
-2011-07-13 -00031955109
XMM6 2011-07-03 0678380501 XMM-Newton/Epic-pn 19
XRT8 2011-07-14 00031955110 Swift/XRT 80
-2011-07-28 -00031955124
XMM7 2011-07-15 0678380601 XMM-Newton/Epic-pn 30
XMM8 2011-07-15 0678380701 XMM-Newton/Epic-pn 19
XRT9 2011-07-29 00031955125 Swift/XRT 77
-2011-08-12 -00031955139
XRT10 2011-08-13 00031955140 Swift/XRT 56
-2011-08-27 -00031955154
XMM9 2011-08-14 0678380801 XMM-Newton/Epic-pn 29
XMM10 2011-08-27 0678380901 XMM-Newton/Epic-pn 30
XRT11 2011-08-28 00031955155 Swift/XRT 108
-2011-10-01 -00031955190
XMM11 2011-09-06 0678381001 XMM-Newton/Epic-pn 28
XMM12 2011-09-06 0678381101 XMM-Newton/Epic-pn 28
XMM13 2011-10-02 0678381201 XMM-Newton/Epic-pn 28
XRT12 2011-10-03 00031955191 Swift/XRT 110
-2011-11-11 -00031955230
XRT13 2011-11-12 00031955231 Swift/XRT 132
-2011-12-21 -00032200015
XRT14 2011-12-22 00032200016 Swift/XRT 104
-2012-01-30 -00032200055
XRT15 2012-01-31 00032200056 Swift/XRT 116
-2012-03-17 -00032200100
XRT16 2012-03-18 00032200101 Swift/XRT 120
-2012-06-07 -00032200175
XRT17 2012-06-08 00032200101 Swift/XRT 156
-2012-09-30 -00032526046
III Results
III.1 Light curves
The long-term X-ray light curve obtained from the Swift /XRT showed a high degree of variation during the TDE of Swift J1644+57. The source counts changed dramatically by orders of magnitude (see Fig. 1) within the span of $\sim 600$ days. We split the energy band of Swift into 0.31.50.3-1.5 keV & 1.510.01.5-10.0 keV ranges. As evident from Fig. 1 and Fig. 2, the hard band dominated the entire episode with sudden dips occasionally observed (Mangano et al., 2016).
We performed cross-correlation analysis using $\zeta$-discrete cross-correlation function (ZDCF [^3] ZDCF: http://www.weizmann.ac.il/particle/tal/research-activities/software, (Alexander, 1997)) on the long-term light curve obtained from the Swift /XRT satellite. $\zeta$-DCF method provides a better estimate of the cross correlation function (CCF) for unevenly sampled data sets and considers the non-linearity while interpolating the lightcures. The cross-correlation study yields an optimal correlation of the 0.3-1.5 keV and 1.5-10 keV energy bands, having a correlation coefficient $\rho_{max}=0.95$. The correlation pattern peaked at a negative delay of $1.09\pm 5.9$ days. Considering the measurement error, this could be considered as a zero-delay between the two energy bands of the X-ray. A similar zero-delay was also observed for an outbursting GBH, named XTE J1550-564, where a relatively smaller size of the accretion disk is attributed to the absence of any delay (Chatterjee et al., 2020). The light curve, along with the $\zeta$-discrete cross-correlation are presented in Fig. 1.
Figure 2: Left Panel: Hardness ratio of Swift J1644+57 is presented. The source exhibited hard spectra throughout the tidal disruption event. The colorbar on top represents the MJD from the start of the event. Right Panel: Correlated variability between 0.2–2 keV and 3–10 keV is observed from XMM-Newton data. The Pearson correlation coefficient is 0.96.
From the light curve of Swift /XRT, we obtained the Hardness-Ratio (HR) as presented in Mangano et al. (2016) and Seifina et al. (2017). We found that the source counts in the hard band were always greater (see Fig. 2) than in the soft band. As the event progressed, the contribution of the hard band in the light curve also increased making harder spectra at the late time of the tidal disruption event. The left panel of Fig. 2 shows the ‘HR’-diagram obtained using the first 90 days of data.
Using XMM-Newton data, we calculated the fractional variability ($F_{\rm var}$) (Edelson et al., 1996, 2001; Edelson and Malkan, 2012; Nandra et al., 1997; Vaughan et al., 2003) in the 0.2–2 and 3–10 keV bands. The light curves (presented as $x_{i}$ counts $s^{-1}$) contain uncertainties $\sigma_{i}$ of length $N$ with a mean $\mu$ and standard deviation $\sigma$, related by,
$F_{\rm var}=\sqrt{\frac{\sigma^{2}_{\rm XS}}{\mu^{2}}},$ (2)
where, $\sigma^{2}_{\rm XS}$ is the excess variance (Nandra et al., 1997; Edelson et al., 2002). $\sigma^{2}_{\rm XS}$ is defined as,
$\sigma^{2}_{\rm XS}=\sigma^{2}-\sigma^{2}_{\rm err},$ (3)
where, $\sigma^{2}_{\rm err}$ is the mean squared error. The $\sigma^{2}_{\rm err}$ is given by,
$\sigma^{2}_{\rm err}=\frac{1}{N}\sum_{i=1}^{N}\sigma^{2}_{\rm i}.$ (4)
The normalized excess variance is given by,
$\sigma^{2}_{\rm NXS}=\frac{\sigma^{2}_{\rm XS}}{\mu^{2}}.$ (5)
The uncertainties in $F_{\rm var}$ and $\sigma_{\rm NXS}$ (Vaughan et al., 2003) are given by,
${\rm err}(F_{\rm var})=\sqrt{\left(\sqrt{\frac{1}{2N}}\frac{\sigma_{\rm err}^{2}}{\mu^{2}F_{\rm var}}\right)^{2}+\left(\frac{1}{\mu}\sqrt{\frac{\sigma^{2}_{\rm err}}{N}}\right)^{2}},$ (6)
and
${\rm err}(\sigma^{2}_{\rm NXS})=\sqrt{\left(\sqrt{\frac{2}{N}}\frac{\sigma_{\rm err}^{2}}{\mu^{2}}\right)^{2}+\left(\sqrt{\frac{\sigma_{\rm err}^{2}}{N}}\frac{2F_{\rm var}}{\mu}\right)^{2}}.$ (7)
Table 2 presents the results of our variability study in soft and hard bands. The excess variance or $\sigma^{2}_{NXS}$ yielded a negative value, which resulted in an imaginary fractional variability ($F_{var}$) for the XMM5 observation. The maximum of $\sigma^{2}_{NXS}$ or $F_{var}$ is observed during XMM3. The $F_{var}$ of the hard band is found to be higher than its softer counterpart for all XMM-Newton observations. We find the $F_{var}$ between 0.2-2 keV and 3-10 keV, presented in the right panel of Fig. 2, are highly correlated with a correlation coefficient of 0.96.
Table 2: Variability statistics in various energy ranges are shown in this table. We have opted for 100 sec time bins for variability analysis. In some cases, the average error of observational data exceeds the limit of $1\sigma$, resulting in negative excess variance. In such cases, we have imaginary $F_{var}$, which are not shown in the table.
ID Energy band NN $x_{max}$ $x_{min}$ $\frac{x_{max}}{x_{min}}$ $\sigma^{2}_{NXS}$ $F_{var}$
keV Count/s Count/s $(10^{-2})$ $(\%)$
XMM1 0.2-2.0 451 8.34 3.69 2.26 $21\pm 0.08 11.$$1.0\pm 0.51$
XMM2 0.2-2.0 378 6.39 1.77 3.56 $1.59\pm 0.16$ $12.6\pm 0.78$
XMM3 0.2-2.0 301 1.46 0.02 60.8 $27.4\pm 2.45$ $52.4\pm 2.83$
XMM4 0.2-2.0 540 1.72 0.37 4.67 $2.92\pm 3.19$ $17.8\pm 1.04$
XMM5 0.2-2.0 302 0.30 0.03 11.8 $0.04\pm 2.95 -$-
XMM6 0.2-2.0 491 1.52 0.18 8.60 $3.90\pm 0.50$ $19.7\pm 1.41$
XMM7 0.2-2.0 322 1.01 0.13 8.00 $2.06\pm 0.52$ $14.3\pm 1.92$
XMM8 0.2-2.0 544 0.94 0.15 6.16 $2.12\pm 0.41$ $15.6\pm 1.46$
XMM9 0.2-2.0 513 0.36 0.03 14.3 $0.51\pm 0.14$ $7.2\pm 9.80$
XMM10 0.2-2.0 521 0.49 0.02 20.1 $27\pm 0.84 11.$$1.2\pm 3.75$
XMM11 0.2-2.0 386 0.43 0.02 17.3 $0.63\pm 1.41$ $7.9\pm 8.89$
XMM12 0.2-2.0 389 0.77 0.04 17.8 $1.12\pm 0.91$ $10.3\pm 4.41$
XMM1 3.0-10.0 451 6.57 2.25 2.93 $2.01\pm 0.12$ $14.1\pm 0.63$
XMM2 3.0-10.0 378 5.28 3.33 2.92 $3.10\pm 0.22$ $17.6\pm 0.89$
XMM3 3.0-10.0 300 1.62 0.02 66.6 $44.7\pm 3.53$ $69.0\pm 3.80$
XMM4 3.0-10.0 540 2.04 0.19 10.8 $8.15\pm 0.47$ $28.4\pm 1.20$
XMM5 3.0-10.0 299 0.26 0.03 10.1 $0.03\pm 0.03 -$-
XMM6 3.0-10.0 491 2.34 0.20 11.4 $10.4\pm 0.63$ $32.3\pm 1.42$
XMM7 3.0-10.0 322 1.36 0.26 5.31 $4.81\pm 0.59$ $21.8\pm 1.60$
XMM8 3.0-10.0 544 1.53 0.23 6.54 $4.32\pm 0.41$ $20.7\pm 1.17$
XMM9 3.0-10.0 513 0.36 0.03 14.0 $30\pm 1.21 11.$$1.2\pm 5.41$
XMM10 3.0-10.0 521 0.68 0.03 27.1 $4.61\pm 0.86$ $21.5\pm 2.10$
XMM11 3.0-10.0 389 0.44 0.03 17.3 $2.54\pm 1.38$ $15.9\pm 4.40$
XMM12 3.0-10.0 387 0.92 0.05 17.7 $7.51\pm 1.05$ $27.4\pm 2.15$
III.2 Spectra
To understand the X-ray spectral properties of Swift J1644+57, we used the phenomenological model along with the physical model. During fitting, we employed stacked XRT spectra to obtain a broad idea of spectral variations. All errors are quoted with 90% confidence level, unless specified otherwise.
We started the spectral fitting with powerlaw model, and the model in XSPEC reads as: TBabs*zTBabs*(powerlaw).
The galactic hydrogen column density ($N_{H}$) was kept frozen at $0.016\times 10^{22}\mathrm{cm}^{-2}$ throughout the entire spectral fitting process. The initial fits with the powerlaw model provide a good fit with $\chi^{2}_{red}\sim 1$. Table 3 presents the findings of the fits. During the first observation, which spans from MJD 55651 to MJD 55665, the spectral index of the source was observed to be $1.79\pm 0.02$. Corresponding $N_{H}$ was found to be $1.75\pm 0.04\times 10^{22}\mathrm{cm}^{-2}$. The normalization of the powerlaw component was observed to be the highest $0.0177\pm 0.005$ ph cm-2 s-1 among all other observations. Our second observation was made by stacking Swift /XRT spectra within the duration of MJD 55666 to MJD 55681. The spectrum was fitted with a power law index $\Gamma=1.62\pm 0.01$ and corresponding $N_{H}$ was found to be $1.44\pm 0.03\times 10^{22}\mathrm{cm}^{-2}$. The normalization and flux, presented in Table 3, were found to be decreased compared to the first observation. This is consistent with the light curve. We found the highest reduced chi-square ($\chi^{2}_{red}\sim 1.21$) for this particular observation. The fitted unfolded spectrum is presented in the left panel of Fig. 3. The following observation considered the stacked spectra within MJD 55682 to MJD 55694. The source flux decreased substantially in the XRT3 observation. We found a higher $N_{H}$ compared to the previous observation. The spectral indices had increased in XRT4 and XRT5 compared to the previous observations, XRT2 and XRT3. Apart from that, the $N_{H}$ had increased, and the corresponding flux decreased. We found a general trend of decreasing $\Gamma$ up to XRT11. The corresponding $N_{H}$ also increased, and the luminosity exhibited a decreasing pattern. The findings were comparable with the findings of Seifina et al. (2017). However, from XRT12, the spectral index started to increase. However, it should be noted that the errors corresponding to the spectral indices became larger as the source became dim and data quality was reduced. For XRT16 observation, we found the source spectrum showed a steeper $\Gamma$ at $1.72\pm 0.22$, comparable to the initial days of the TDEs. The left Panel of Fig. 4 shows the $90\%$ confidence contours of $N_{H}$ and $\Gamma$. The spectral parameter variations, fitted with the powerlaw model, are presented in Fig. 5.
Figure 3: left to right: XRT2, XRT6, and XRT10 fitted unfolded spectra are presented for various models. The red line represents the Optxagnf, green represents the Nthcomp, and orange represents the powerlaw model. Corresponding residues are plotted in the bottom panels.
XMM-Newton observed Swift J1644+57 twelve times within the duration of the initial phase of the tidal disruption event ranging from MJD 55667 to MJD 55836. We fitted the XMM-Newton data with the powerlaw model to compare the parameter variations with the stacked observations created using Swift data. During the first observation, referred to as XMM1, we found that the source spectrum had a spectral index of $1.69\pm 0.01$ with a relatively lower hydrogen column density of $1.29\pm 0.01\times 10^{22}\mathrm{cm}^{-2}$. The powerlaw normalization and flux are found to be highest (see Table 3) in that particular observation. In the following observation, XMM2, the spectrum hardened with $\Gamma=1.54\pm 0.01$ and associated $N_{H}$ increased to $1.42\pm 0.02\times 10^{22}\mathrm{cm}^{-2}$. The normalization and the flux were reduced substantially. Later, on MJD 55697, we fitted the XMM3 spectrum where $\Gamma$ was found highest among all the XMM-Newton observations, having a value of $1.87\pm 0.05$. The flux decreased dramatically to $11.14\times 10^{-11}$ from $10.15\times 10^{-11}$ erg cm-2 s-1 . This sudden dip in the flux could be associated with the dips found in the long-term light curve (Mangano et al., 2016; Seifina et al., 2017). During XMM4 observation, the source spectrum returned to a harder spectrum with $\Gamma=1.55\pm 0.02$. The concurrent flux increased from the previous observation. Later, in the XMM5 observation, we found an upward variation of the spectral index with $\Gamma=1.59\pm 0.08$. However, considering the error, the spectral index could be in a similar range as the XMM4 observation. The normalization and flux decreased rapidly, and the observed values are the lowest among all the XMM-Newton observations. Thereafter, both the flux and normalization increased with decreased spectral index. Post XMM5, the powerlaw fitted spectra exhibited a steady decrease in the flux with occasional variation in the spectral index. The source normalization also varied from $13.56\times 10^{-3}$ to $0.78\times 10^{-3}$ ph cm-2 s-1 throughout the XMM-Newton observations.
[…]
III.2.1 Optxagnf
The Optxagnf model, proposed by Done et al. (2012), provides an in-depth understanding of the spectral parameters as the parameters are generated from a physical picture of the accretion disk around the AGNs. The model-fitted parameters include the luminosity, coronal radius parameter ($R_{cor}$), spin of the central black hole, soft-excess temperature, optical depth, spectral index, and the fraction of Comptonized photons. It should be noted that Optxagnf is the only publicly available model to fit the X-ray spectra of AGNs where $R_{cor}$ could be estimated. We employed this model to fit the XRT spectra. The model in xspec reads as: TBabs*zTBabs*(optxagnf).
The model spectrum computes both soft-excess and Comptonized photons. The accretion rate governed by the mass of the central black hole determines the luminosity of the source (in units of Eddington luminosity - $log(L/L_{Edd})$). The geometry of the Optxagnf  model separates the accretion disk into three regions. First, the disk emission is produced through a color temperature-corrected blackbody emission from radii $R_{\rm out}>r>R_{\rm cor}$, where $R_{\rm out}$ and $R_{\rm cor}$ are the outer edge of the disk and the corona, respectively. Luminosity is computed keeping the distance and mass constant, while the mass-to-energy conversion efficiency factor ($\eta$) depends on the spin. Second, at $r<R_{\rm cor}$, the emission is dominated by Comptonized emission from a warm and optically thick medium. The third region where inverse Comptonization occurs is the optically thin corona around the disk, and the fraction of inverse Comptonization can be estimated from the $f_{pl}$ parameter during the spectral fitting. The temperature and optical depth of the warm, optically thick medium could be found from $kT_{s}$ and $\tau$, respectively. The model calculates the primary emission of Comptonized photons using the Nthcomp model, where the seed photon temperature is kept fixed at $r=R_{\rm cor}$, and the electron temperature of the optically thin hot corona is frozen at 100 keV. We freeze the mass of the black hole at $3\times 10^{6}~M_{\odot}$ (Bloom et al., 2011b). The Galactic $N_{H}$ along the line of Swift J1644+57  remained frozen at $0.016\times 10^{22}\mathrm{cm}^{-2}$ while performing the fitting. According to the model, it was suggested to keep the normalization frozen at unity. The redshift $z=0.354$ and distance $d=1402.8$ Mpc are also kept fixed (Levan et al., 2011). The $log(r_{out})$ is kept fixed at 5 since the apogee of extremely elongated TDEs, capable of explaining lengthy disruption events (Hayasaki et al., 2013; Dai et al., 2015; Hayasaki et al., 2018), could reach up to $10^{4}-10^{5}\,r_{\rm g}$. The remaining parameters are left free to vary. It should be noted that the reported luminosities in Table 4 are isotropic and model-dependent without employing the beaming corrections. From Table 3, we can see that the power law model fits the spectra of Swift J1644+57  well, and the soft X-ray excess is absent in the spectra. Thus, we freeze the $kT_{s}$ at 0.1 keV and $\tau$ at a maximum value of 5.0. Initially, we left the $f_{pl}$ thawed. We observed that the $f_{pl}$ values saturated in most cases. Later, better fits were obtained while fitting the spectra with variable $kT_{s}$ and $\tau$ and is presented in Table 4. We observed that the warm corona temperature and optical depth remained high across all observations except XRT3. We found that the $kT_{s}$ gradually increases as the event progresses and remains within $0.59\leq kT_{s}\leq 5.73$. On the other hand, the variation of optical depth remained non-monotonic within the limit $1.25\leq\tau\leq 7.83$. The variation of $f_{pl}$ remained within the range of $0.48\leq f_{pl}\leq 0.93$ during our observation period. For TDEs, this is expected and is conjectured by Chakraborty et al. (2026). Optxagnf fitted spectra for three epochs are presented in Fig. 3, and the variation of fitted parameters with time is plotted in the right panel of Fig. 5 and are presented in the Table 4.
We fitted the 0.3–10 keV stacked Swift spectra for the entire episode of TDE. The spectra fitted alone with the Optxagnf model provided good statistics with $\chi^{2}_{red}\sim 1$. The model-fitted spectral parameters are presented in Table 4. We found the luminosity of the source exceeded nearly 500 times the Eddington luminosity during the initial few days of the event. This agrees with the earlier studies like Zauderer et al. (2013) and is also seen from the current fitting with Optxagnf model. Over time, the luminosity of the source decreased nearly 200 times from the initial phase, but never went into the sub-Eddington regime during the period of our observation. The variation of $N_{H}$ remained similar to what has been observed during the fitting with the powerlaw model. The spin of the central black hole remained within the range of $0.4<a^{*}<0.8$, making it an intermediate spinning black hole. During the initial days, the coronal parameter of the source was observed at $31\pm 2~r_{\rm g}$. The right panel of Fig. 4 represents the confidence contours of spin ($a$) vs $R_{cor}$. The coronal radius parameter was found to vary within $31<R_{cor}<65~r_{\rm g}$. We only used the stacked Swift data for Optxagnf fitting as it provides better data quality to understand the long-timescale variations of Swift J1644+57 during its tidal disruption episode.
Table 3: powerlaw fitting result for the 0.3–10 keV energy range.
ID MJD $N_{H}$ $\Gamma$ N log(Flux) $\chi^{2}/dof$
(days) $(10^{22}~cm^{-2})$ $(10^{-3})$ ($10^{-11}$ erg/cm2/s)
Swift /XRT
XRT1 $55658$ $1.75^{+0.04}_{-0.04}$ $1.79^{+0.02}_{-0.02}$ $17.7^{+5.50}_{-5.50}$ 09.94-09.94 $902.74/827$
XRT2 $55673.5$ $1.44^{+0.03}_{-0.03}$ $1.62^{+0.01}_{-0.01}$ $10.8^{+2.12}_{-2.12}$ 10.09-10.09 $1047.47/861$
XRT3 $55688$ $1.67^{+0.08}_{-0.08}$ $1.63^{+0.04}_{-0.04}$ $4.51^{+2.74}_{-2.74}$ 10.48-10.48 $741.23/866$
XRT4 $55702.5$ $1.90^{+0.08}_{-0.08}$ $1.75^{+0.04}_{-0.04}$ $3.10^{+1.73}_{-1.73}$ 10.63-10.63 $525.19/703$
XRT5 $55717.5$ $1.99^{+0.10}_{-0.10}$ $1.74^{+0.05}_{-0.05}$ $2.20^{+1.60}_{-1.60}$ 10.83-10.83 $574.73/643$
XRT6 $55731.5$ $1.95^{+0.11}_{-0.11}$ $1.61^{+0.05}_{-0.05}$ $1.47^{+1.13}_{-1.13}$ 10.96-10.96 $615.79/640$
XRT7 $55747$ $2.13^{+0.15}_{-0.15}$ $1.62^{+0.07}_{-0.07}$ $0.96^{+0.96}_{-0.96}$ 11.14-11.14 $596.64/594$
XRT8 $55763$ $1.89^{+0.13}_{-0.13}$ $1.45^{+0.07}_{-0.07}$ $1.26^{+0.12}_{-0.12}$ 10.96-10.96 $555.14/596$
XRT9 $55778$ $1.94^{+0.16}_{-0.16}$ $1.56^{+0.07}_{-0.07}$ $1.33^{+0.14}_{-0.14}$ 10.98-10.98 $508.38/544$
XRT10 $55793$ $2.21^{+0.20}_{-0.20}$ $40^{+0.09}_{-0.09} 1.1$$12^{+0.15}_{-0.15}$ 10.98-10.98 $498.03/520$
XRT11 $55818$ $2.04^{+0.22}_{-0.22}$ $1.44^{+0.10}_{-0.10}$ $0.46^{+0.07}_{-0.07}$ 11.39-11.39 $431.81/475$
XRT12 $55856.5$ $2.14^{+0.19}_{-0.19}$ $1.57^{+0.10}_{-0.10}$ $0.72^{+0.09}_{-0.09}$ 11.25-11.25 $466.99/509$
XRT13 $55896.5$ $2.15^{+0.19}_{-0.19}$ $1.50^{+0.09}_{-0.09}$ $0.58^{+0.07}_{-0.07}$ 11.31-11.31 $441.98/519$
XRT14 $55936$ $2.18^{+0.39}_{-0.39}$ $1.61^{+0.16}_{-0.16}$ $0.38^{+0.10}_{-0.10}$ 11.54-11.54 $274.05/309$
XRT15 $55980$ $1.94^{+0.24}_{-0.24}$ $1.55^{+0.12}_{-0.12}$ $0.32^{+0.06}_{-0.06}$ 11.59-11.59 $337.79/371$
XRT16 $56044.5$ $2.23^{+0.55}_{-0.55}$ $1.72^{+0.22}_{-0.22}$ $0.21^{+0.08}_{-0.08}$ 11.84-11.84 $166.80/197$
XRT17 $56143$ $1.76^{+0.80}_{-0.80}$ $1.28^{+0.34}_{-0.34}$ $0.04^{+0.02}_{-0.02}$ 12.40-12.40 $99.32/104$
XMM-Newton
XMM1 $55667$ $1.29^{+0.01}_{-0.01}$ $1.69^{+0.01}_{-0.01}$ $13.56^{+0.10}_{-0.10}$ 10.02-10.02 $224.69/174$
XMM2 $55681$ $1.42^{+0.02}_{-0.02}$ $1.54^{+0.01}_{-0.01}$ $8.87^{+0.10}_{-0.10}$ 10.15-10.15 $186.68/173$
XMM3 $55697$ $1.33^{+0.06}_{-0.06}$ $87^{+0.05}_{-0.05} 1.1$$8^{+0.07}_{-0.08} 11$.14-11.14 $141.20/133$
XMM4 $55711$ $1.49^{+0.04}_{-0.03}$ $1.55^{+0.02}_{-0.02}$ $2.48^{+0.07}_{-0.07}$ 10.71-10.71 $193.00/167$
XMM5 $55745$ $1.41^{+0.14}_{-0.13}$ $1.59^{+0.08}_{-0.08}$ $0.27^{+0.03}_{-0.03}$ 11.68-11.68 $97.00/90$
XMM6 $55757$ $1.61^{+0.05}_{-0.05}$ $1.39^{+0.02}_{-0.02}$ $1.68^{+0.06}_{-0.06}$ 10.80-10.80 $186.57/162$
XMM7 $55769$ $1.62^{+0.06}_{-0.06}$ $1.41^{+0.03}_{-0.03}$ $1.44^{+0.07}_{-0.06}$ 10.87-10.87 $190.57/151$
XMM8 $55787$ $1.63^{+0.05}_{-0.05}$ $1.34^{+0.02}_{-0.02}$ $1.41^{+0.05}_{-0.05}$ 10.85-10.85 $179.59/163$
XMM9 $55800$ $1.43^{+0.10}_{-0.10}$ $1.31^{+0.05}_{-0.05}$ $0.30^{+0.02}_{-0.02}$ 11.51-11.51 $107.40/129$
XMM10 $55810$ $1.58^{+0.08}_{-0.08}$ $1.39^{+0.04}_{-0.04}$ $0.58^{+0.03}_{-0.03}$ 11.26-11.26 $154.57/145$
XMM11 $55824$ $1.51^{+0.11}_{-0.12}$ $1.41^{+0.06}_{-0.06}$ $0.37^{+0.03}_{-0.03}$ 11.46-11.46 $154.39/121$
XMM12 $55836$ $1.51^{+0.08}_{-0.08}$ $1.34^{+0.04}_{-0.04}$ $0.78^{+0.04}_{-0.05}$ 11.10-11.10 $167.89/144$
Figure 4: Contours of confidence are plotted for Swift /XRT observations fitted with powerlaw and Optxagnf model.
IV Discussion
We studied Swift J1644+57 during its tidal disruption event of 2011-2012 using Swift and XMM-Newton data within the energy range 0.3-10 keV. We employed the powerlaw as our phenomenological model and a physical model, such as Optxagnf, to understand the behavior of the source as the event progresses. Being a jetted-TDE, the source exhibited variability in both the spectral and temporal domains and is one of the most studied tidal disruption events to date. During our analysis, we found that there are several similarities between TDEs and the declining phase of an outbursting GBH. The similarities between GBH transients and TDEs are seen from their spectral energy distributions and the decaying luminosity with time. The X-ray spectral indices of GBHs are found to decrease in the luminosity decay phase (McClintock et al., 2009). Likewise, the decreasing tendency of the spectral index has also been observed in Swift J1644+57 (Seifina et al., 2017). Recent NICER observations (Gendreau et al., 2012) of MAXI J1820+070 provided evidence of a dynamically contracting corona for the onset stage of the flare (Kara et al., 2019). Jana et al. (2016) explained, using the two-component advective flow model (Chakrabarti and Titarchuk, 1995), why the Compton cloud expanded at the declining phase of a low inclination GBH, MAXI J1836-194, during the 2011 outburst. The variation of the corona size is also observed for AGN IRAS 13224-3809 (Alston et al., 2020). Apart from $R_{cor}$ variation, relativistic jet launching is another phenomenon that GBH transients and TDEs have in common (Mirabel and Rodríguez, 1994; Patra et al., 2019; Espinasse et al., 2020; Blandford et al., 2019; Zauderer et al., 2011).
IV.1 Nuclear Properties
X-ray emission from Swift J1644+57 has been observed for around 600 days. The hard X-ray component was found to dominate, and the radiation is conjectured to be beamed (Zauderer et al., 2013) and many authors. From our spectral analysis, we find that the source exhibited variations in hydrogen column density ($N_{H}$). Fig. 5, represents the variation of powerlaw fitted parameters. We find that the hydrogen column density, at the beginning, was higher at $1.75\pm 0.04\times 10^{22}\mathrm{cm}^{-2}$. Later, in XRT2, the $N_{H}$ dropped to $1.44\pm 0.04\times 10^{22}\mathrm{cm}^{-2}$. Post XRT2, the spectra showed a gradual increase in the value of $N_{H}$ and plateaued at a late stage. Similar observational findings are also seen through XMM-Newton spectra fitted with powerlaw model and Swift spectra fitted with the Optxagnf model. The variation of $N_{H}$ could be a generic property of the jetted-TDEs.
Seifina et al. (2017) reported that the spectral index of individual spectra observed by Swift /XRT saturates after MJD 55730; i.e., after eighty days since the start of the event. We used stacked spectra of Swift /XRT where the spectral indices are observed to be decreasing up to MJD 55800. However, after that, the spectral indices increased. But, finally, at XRT17 observation, the spectral index was $1.28\pm 0.34$. Both Swift and XMM-Newton data are employed to measure the spectral indices variation within MJD 55650 to MJD 55800, where a similar pattern has been found. Using the Optxagnf model, we have also observed similar variations. The increase in the X-ray spectral indices after 150 days could be associated with the increased optical activity after $\sim 30-50$ days since the initial observations (Levan et al., 2016). We find that $N_{H}$ and $\Gamma$ are weakly anti-correlated with a Pearson Correlation coefficient of -0.22.
The Optxagnf model provides several parameters of the source. Among them, we find that the luminosity, in terms of $\log(L/L_{Edd})$, is highly sensitive to the fitting. The fitted spectra exhibited a variation of the order of 2 with the range of our observations. During the initial days, Swift J1644+57 had a bolometric luminosity of around 500 times the Eddington limit. In the entire observation, the source luminosity was in the super-Eddington regime. The luminosity followed a steady decrease from the initial days, with occasional variation in between, which can be associated with the intrinsic photon count variation in the light curve. During XRT17, the last observation, the source luminosity decreased to nearly 3 times the Eddington limit. We find the source luminosity weakly correlated with the hydrogen column density, with a Pearson Correlation coefficient of -0.31.
The spin of Swift J1644+57, measured using the $a^{*}$ parameter of Optxagnf model, exhibited a region of an intermediate spinning black hole. We find $a^{*}$ to vary from 0.4 to 0.8. The value of the spin aligns with the earlier claim by Tchekhovskoy et al. (2014), where they predicted the spin of the source should be greater than $\geq 0.5$.
Figure 5: Variation of spectral parameters of Swift J1644+57 are presented for powerlaw and optxagnf model. Only stacked Swift/XRT data are employed to fit optxagnf model.
IV.2 Coronal Variations
The hard X-ray photons, emerging from the vicinity of a black hole, are thought to originate from a corona which can be formed through magnetic reconnection in the accretion disk (Galeev et al., 1979; Liu et al., 2002). There are a few practical models of corona that have been proposed so far: Lampost model (Martocchia and Matt, 1996; Miniutti and Fabian, 2004) provides physical explanations of the X-ray spectra from the AGNs; Done et al. (2012) constructs Optxagnf model, computes the size of the corona, $R_{cor}$ , from the spectral fitting. Also, it has been observed that the nature of the corona is dynamic for GBHs (Kara et al., 2019) and for the AGNs (Alston et al., 2020). That is why it is crucial to explore how the corona size evolves with time for TDEs.
Figure 6: Optxagnf fitted spectral parameters variation with each other of Swift J1644+57 are presented. Panel (a) shows the variation of $N_{H}$ with respect to the normalized luminosity. The variation of $N_{H}$ with respect to $\Gamma$ is shown in panel (b). Panel (c) demonstrates the weak correlation between $\Gamma$ and scattering fraction $f_{pl}$. Variation of the $\Gamma$ and $\tau$ can be observed in panel (d). Panel (e) presents the variation of $R_{cor}$ and spin parameter. Uncorrelated variation between spin and $R_{cor}$ can be observed in panel (e).
To measure the $R_{cor}$ variation of Swift J1644+57, we employ the Optxagnf model to determine the spectral fits using the stacked Swift /XRT data. Through our spectral fitting, we find that at the peak luminosity of Swift J1644+57, the radial size of the corona was $31\pm 2.3~r_{\rm g}$. Subsequently, an almost similar size of the $R_{cor}$ was found in the XRT2 and XRT3 observational epochs. The size of the corona increases $35\pm 3.6~r_{\rm g}$ in the XRT4 epoch. Later on, in the XRT5 and XRT6 epochs, the $R_{cor}$ keeps increasing to $43\pm 3.7~r_{\rm g}$ and $51\pm 4.3~r_{\rm g}$, respectively. During the XRT7 observation, the $R_{cor}$ is $58\pm 4.1~r_{\rm g}$. Up to the XRT7 observation, we observed a sharp rise in the $R_{cor}$ . The fitted value of the $R_{cor}$ is $61\pm 4.2~r_{\rm g}$ at XRT8 and drops to $58\pm 4.5~r_{\rm g}$ at XRT10. From the XRT10 to XRT17 observations, the $R_{cor}$ varies little from $58\,r_{\rm g}$ to $64\,r_{\rm g}$, indicating a saturation level of the parameter. All estimated values of the coronal parameters are depicted in Fig. 8. In addition, we find through our analysis that the $R_{cor}$ is anti-correlated with the luminosity $\log(L/L_{\rm Edd})$ with the Pearson correlation coefficient: 0.87-0.87.
Recently, Mummery and Balbus (2021) proposed the disk-corona model to explain non-thermal hard-X-ray emissions from X-ray TDEs. According to their prescription, the size of the corona would depend on the disk luminosity as follows:
[…]
where
$f(l)=\frac{l_{I}-l(t)}{l_{I}-l_{C}},$
$l_{I}$, $l_{C}$, and $l(t)$ are the initial, final, and the time-dependent Eddington ratios, respectively. Starting with an initial $R_{cor}$ $R_{I}$, the corona expands following the decay pattern of the luminosity profile and finally saturates to a value of $R_{C}$. We adopt $R_{I}=30\,r_{\rm g}$ and $R_{C}=65\,r_{\rm g}$. The normalized disk luminosity, $\log(L/L_{\rm Edd})$, at each observational epoch is displayed in Table 4. The predicted $R_{cor}$ at each epoch using Equation (8) is denoted with red circles in Fig. 8. The error in the theoretical values is a contribution arising from the observational error in the measurement of luminosity as presented in Table 4. We find that the evolution of theoretically speculated corona size agrees with that of the observational $R_{cor}$ parameter in Swift J1644+57, and considering the error bars, their variation remains similar.
Fig. 9 gives a schematic view of the disk-corona-jet system for the early and late stages of Swift J1644+57. The physical scale of the disk and corona fits within the range of $2R_{\rm t}\sim 100\,r_{\rm g}$ for a $10^{6}~M_{\odot}$ SMBH, where $R_{\rm t}$ is given by Equation (1). We presume that, in the initial days, the accretion rate was at the super-Eddington rate, and the corona was small. As the days progressed, the accretion rate reduced, and the magnetic flux became stronger, as suggested by Tchekhovskoy et al. (2014). This allows the disk to have a bigger corona than that of the initial state. Although the disk is yet in the super-Eddington regime at this moment, the bigger corona is likely to produce a larger number of high-energy X-ray photons via inverse Compton scattering. Within the first hundred days, the jet flux increases monotonically and is anti-correlated with the X-ray flux (Berger et al., 2012; Seifina et al., 2017), suggesting that the jet region also expands during this period. Depending on the quality and amount of the X-ray observational data, one can measure the coronal variation of the other three jetted TDEs and non-jetted TDEs using the same analysis as ours. This is also useful for improving the theoretical corona model in TDEs. More recently, Chakraborty et al. (2026) argued that the accretion disk of X-ray bright TDEs could be composed of a variable corona having lower electron temperature and higher optical depth, and can also be observed from the variation of parameters presented in Table 4.
Figure 7: Markov Chain Monte-Carlo plot for parameter variations of XRT2 and XRT15 are presented. Goodman-Weare (EMCEE) (Foreman-Mackey et al., 2013) is opted for 32 walkers and a chain length of 1 million iterations, where the transient phase is removed by burning the chain up to 9984 iterations. The spin parameter was kept frozen at $a^{*}=0.60$ to perform the MCMCs.
Figure 8: Coronal variation with time is presented for Swift J1644+57. Swift /XRT fitted values are presented by green-diamond points. The red circles denote the theoretical values of $R_{cor}$ as predicted by Mummery and Balbus (2021).
Figure 9: Edge-on view schematic diagram of $R_{cor}$ variation in the early and late stages of Swift J1644+57. The system comprises a disk around the SMBH, a corona on the disk, and a jet. The disk size is twice the tidal disruption radii, which is approximately estimated to be $2R_{\rm t}\sim 100\,r_{\rm g}$ for the $10^{6}M_{\odot}$ SMBH and a solar-type star (see equation 1).
IV.2.1 Viability of TDE Corona Models
It should be noted that, to date, there exists no publicly available model in XSPEC to fit the hyper accretion regime of Swift J1644+57. The source was one of the very few jetted TDEs ever observed with a QPO feature. The initial accretion rate was $\sim 1000~L_{Edd}$ and was widely explored for its unique attributes. While choosing the model, one must be extremely cautious. Optxagnf is the only model in XSPEC that applies to the super-Eddington accretion regime of the AGNs, and measures the coronal radius parameter.
[…]
Recently, Chakraborty et al. (2026) has proposed a coronal emission for TDEs (see their Fig. 7) where the optical depth could be higher, accompanied by a lower electron temperature. The parameters obtained from the Optxagnf model, presented in Table 4, exhibit a similar nature of corona. Although the morphology of the TDE corona could depart from the structure of the warm corona, the thermal properties, such as high $\tau$ and $kT_{s}$, resemble the warm corona of the Optxagnf model. The variation of $R_{cor}$ could be observed from the MCMC plots presented in the Fig. 7. However, it should be noted that Optxagnf, developed for AGNs, may not be physically self-consistent for hyper-Eddington jetted TDEs, such as for Swift J1644+57.
The high optical depth ($\tau$) & low ($kT_{s}$) corona could form in the case of radiation pressure supported tori or super-Eddington thick disks, which were investigated by Abramowicz et al. (1978); Kozlowski et al. (1978); Abramowicz et al. (1980); Paczyńsky and Wiita (1980); Sikora (1981); Fukue (1982); Chakrabarti (1985b); Madau (1988); Ulmer (1997); Jiang et al. (2019); Jiang and Dai (2024). Thick disks are capable of launching collimated jets (Rees et al., 1982; Chakrabarti, 1985a) and could also produce QPOs through oscillations (Jaroszynski, 1986). The radiative properties of the TDEs need to be explored in this regime. However, there are no models in XSPEC with the thick disk feature. Furthermore, under the ideal situation, one could be exploring the Magnetically Arrested Disks (MAD, Narayan et al. (2003); Tchekhovskoy et al. (2014)) to fit the broadband SED spectra of the jetted TDEs. In future work, we will incorporate a simplified thick disk model in XSPEC to examine the parameters for several X-ray bright TDEs.
Table 4: Optxagnf fitted spectral parameter are presented. We used stacked Swift/XRT over XMM-Newton as the former provides more degrees of freedom over the latter. The following parameters are kept frozen while fitting: $N_{H}^{Gal}=0.016\times 10^{22}\mathrm{cm}^{-2}$, model normalization $N_{optx}=1$, redshift $z=0.354$, distance $d=1402.8$ Mpc, and $log(r_{out})=5$.
ID MJD $N_{H}$ $\log(\frac{L}{L_{Edd}})$ $a^{*}$ $R_{cor}$ $kT_{s}$ $\tau$ $\Gamma$ $f_{pl}$ $dof/chi^{2}$
(days) ($10^{22}cm^{-2}$) ($r_{g}$) (keV)
XRT1 55658 $2.14^{+0.06}_{-0.07}$ $2.69^{+0.13}_{-0.12}$ $0.62^{+0.03}_{-0.04}$ $31^{+2.1}_{-2.3}$ $1.29^{+0.34}_{-0.33}$ $4.99^{+0.23}_{-0.24}$ $1.70^{+0.04}_{-0.04}$ $0.87^{+0.04}_{-0.07}$ $887/830$
XRT2 55673.5 $1.58^{+0.12}_{-0.14}$ $2.55^{+0.21}_{-0.20}$ $0.58^{+0.11}_{-0.13}$ $31^{+3.3}_{-4.1}$ $4.41^{+1.29}_{-1.28}$ $4.99^{+0.31}_{-0.27}$ $1.62^{+0.09}_{-0.10}$ $0.88^{+0.04}_{-0.04}$ $1058/938$
XRT3 55688 $2.41^{+0.04}_{-0.03}$ $2.42^{+0.11}_{-0.15}$ $0.49^{+0.04}_{-0.06}$ $32^{+2.5}_{-2.2}$ $0.59^{+0.30}_{-0.25}$ $2.94^{+0.22}_{-0.22}$ $1.79^{+0.07}_{-0.07}$ $0.48^{+0.03}_{-0.03}$ $722/689$
XRT4 55702.5 $1.95^{+0.08}_{-0.07}$ $1.88^{+0.13}_{-0.13}$ $0.60^{+0.05}_{-0.04}$ $35^{+3.1}_{-3.6}$ $2.15^{+0.29}_{-0.26}$ $1.25^{+0.15}_{-0.15}$ $1.77^{+0.08}_{-0.07}$ $0.89^{+0.09}_{-0.08}$ $625/703$
XRT5 55717.5 $2.06^{+0.05}_{-0.05}$ $1.77^{+0.16}_{-0.17}$ $0.63^{+0.06}_{-0.06}$ $43^{+3.7}_{-3.4}$ $1.03^{+0.25}_{-0.22}$ $2.99^{+0.23}_{-0.25}$ $1.76^{+0.06}_{-0.08}$ $0.81^{+0.06}_{-0.08}$ $575/646$
XRT6 55731.5 $2.07^{+0.09}_{-0.08}$ $1.64^{+0.10}_{-0.10}$ $0.61^{+0.07}_{-0.06}$ $51^{+4.1}_{-4.3}$ $1.28^{+0.24}_{-0.26}$ $3.00^{+0.28}_{-0.28}$ $1.69^{+0.10}_{-0.11}$ $0.93^{+0.06}_{-0.06}$ $683/638$
XRT7 55747 $2.15^{+0.11}_{-0.12}$ $1.51^{+0.16}_{-0.15}$ $0.67^{+0.10}_{-0.10}$ $58^{+3.8}_{-4.1}$ $2.22^{+0.31}_{-0.32}$ $1.30^{+0.27}_{-0.27}$ $1.64^{+0.12}_{-0.11}$ $0.90^{+0.03}_{-0.02}$ $593/587$
XRT8 55763 $1.86^{+0.08}_{-0.06}$ $1.88^{+0.13}_{-0.12}$ $0.62^{+0.07}_{-0.08}$ $61^{+4.6}_{-4.1}$ $1.15^{+0.52}_{-0.54}$ $3.46^{+0.35}_{-0.36}$ $1.49^{+0.16}_{-0.16}$ $0.93^{+0.05}_{-0.06}$ $550/594$
XRT9 55778 $1.95^{+0.05}_{-0.09}$ $1.55^{+0.14}_{-0.13}$ $0.54^{+0.04}_{-0.04}$ $58^{+4.5}_{-4.5}$ $3.52^{+0.24}_{-0.24}$ $6.74^{+0.41}_{-0.38}$ $1.45^{+0.09}_{-0.08}$ $0.56^{+0.15}_{-0.14}$ $510/547$
XRT10 55793 $2.03^{+0.13}_{-0.18}$ $1.52^{+0.14}_{-0.14}$ $0.58^{+0.10}_{-0.10}$ $61^{+4.1}_{-4.2}$ $3.13^{+0.33}_{-0.33}$ $7.83^{+0.44}_{-0.39}$ $1.51^{+0.11}_{-0.10}$ $0.55^{+0.17}_{-0.17}$ $511/527$
XRT11 55818 $99^{+0.22}_{-0.13} 1.1$$10^{+0.19}_{-0.19}$ $0.71^{+0.11}_{-0.15}$ $58^{+5.2}_{-5.1}$ $3.88^{+1.23}_{-1.23}$ $6.71^{+0.46}_{-0.46}$ $1.48^{+0.16}_{-0.16}$ $0.65^{+0.21}_{-0.19}$ $426/463$
XRT12 55856.5 $2.23^{+0.13}_{-0.09}$ $0.94^{+0.11}_{-0.12}$ $0.66^{+0.13}_{-0.09}$ $63^{+5.1}_{-5.2}$ $4.30^{+0.65}_{-0.71}$ $5.82^{+0.45}_{-0.47}$ $1.30^{+0.10}_{-0.08}$ $0.81^{+0.07}_{-0.07}$ $444/461$
XRT13 55896.5 $2.18^{+0.20}_{-0.14}$ $1.17^{+0.15}_{-0.13}$ $0.55^{+0.22}_{-0.16}$ $60^{+5.7}_{-6.1}$ $5.42^{+0.60}_{-0.68}$ $5.79^{+0.50}_{-0.51}$ $1.64^{+0.20}_{-0.14}$ $0.64^{+0.15}_{-0.15}$ $440/508$
XRT14 55936 $2.41^{+0.07}_{-0.07}$ $1.07^{+0.16}_{-0.18}$ $0.67^{+0.11}_{-0.12}$ $64^{+5.0}_{-3.4}$ $4.99^{+1.37}_{-1.37}$ $3.52^{+0.17}_{-0.17}$ $1.24^{+0.12}_{-0.12}$ $0.48^{+0.14}_{-0.13}$ $246/300$
XRT15 55980 $2.45^{+0.08}_{-0.11}$ $1.37^{+0.08}_{-0.11}$ $0.53^{+0.14}_{-0.10}$ $61^{+4.3}_{-4.6}$ $5.73^{+0.31}_{-0.33}$ $1.70^{+0.29}_{-0.31}$ $1.46^{+0.08}_{-0.11}$ $0.78^{+0.09}_{-0.09}$ $329/368$
XRT16 56044.5 $2.56^{+0.21}_{-0.17}$ $0.84^{+0.13}_{-0.15}$ $0.76^{+0.12}_{-0.10}$ $65^{+4.2}_{-3.8}$ $2.50^{+0.42}_{-0.44}$ $6.90^{+0.33}_{-0.33}$ $1.35^{+0.10}_{-0.07}$ $0.82^{+0.07}_{-0.06}$ $166/192$
XRT17 56143 $1.92^{+0.20}_{-0.13}$ $0.46^{+0.10}_{-0.10}$ $0.68^{+0.09}_{-0.08}$ $64^{+5.2}_{-4.9}$ $4.96^{+1.30}_{-1.30}$ $5.06^{+0.31}_{-0.32}$ $1.29^{+0.11}_{-0.11}$ $0.76^{+0.08}_{-0.10}$ $99.5/100$
V Summary
We have re-visited the X-ray emission of the jetted-TDE Swift J1644+57 between 2011 and 2012 by analyzing the archival Swift and XMM-Newton data with HEAsoft. Swift J1644+57 showed a massive tidal disruption event during 2011 and 2012, where the flux of the hard (3–10 keV) photons dominated that of the soft (0.2–2 keV) photons over the entire period. We have hypothesized that those hard photons originate from the Compton cloud, i.e., the corona, and have estimated the size of the corona by measuring the size of the hard X-ray emitted region using the publicly available model Optxagnf. Our key findings are summarized as follows.
  1. 1.
    We find that the spectral index non-monotonically decreases with time, demonstrating that the fraction of harder photons increases with time. This indicates that the size of the Compton cloud gets larger because the soft photons are more scattered to be exhausted. We observed that the optical depth ($\tau$) and the temperature of the warm corona ($kT_{s}$) remained high – a result that aligns with the Chakraborty et al. (2026) study.
  2. 2.
    By interpreting the hard X-ray-emitted region as a corona, we have estimated the size of the corona as a function of time. Considering the Optxagnf model, we confirm that the coronal radius parameter $R_{cor}$ increases over time for the observed period. We also find that the evolution of the corona size agrees with the theoretical conjecture made by Mummery and Balbus (2021).
  3. 3.
    We find that the flux of hard and soft photons is highly correlated with the correlation coefficient $\rho_{max}=0.95$ throughout the event. Their correlation pattern has peaked at a zero delay. Moreover, the flux variabilities of the soft (0.2-2 keV) and hard (3-10 keV) photons correlate with each other with a Pearson correlation coefficient of $0.96$. These correlations suggest that the soft and hard band photons are emitted from the same source.
Four jetted TDEs have been observed so far (Cenko et al., 2012; Brown et al., 2015; Andreoni et al., 2022). In the future, more exotic events could be observed by XRISM (Tashiro et al., 2018). The large field of view and imaging capability of the proposed-to-NASA mission AXIS (Mushotzky and AXIS Team, 2019; Reynolds et al., 2023; Safi-Harb et al., 2023; Koss et al., 2025) could provide a detailed structure of the X-ray emitting region of these sources. Also, to perform the short-term timing studies and search for more QPOs among the AGNs, the large effective area and high throughput of the proposed-to-CSA Colibrì (Heyl et al., 2019; Caiazzo et al., 2019) and proposed-to-ESA NewAthena (Cruise et al., 2025) missions will be required in the future. Future proposed X-ray interferometric missions (Cash et al., 2000; Cash, 2003; Gendreau et al., 2004; Chatterjee et al., 2017, 2018; Chatterjee, 2018; Uttley et al.,, 2020; Gandhi, 2023; Weaver et al., 2026) are essential to visualize the coronal variation in the configuration space, much like what EHT (Event Horizon Telescope Collaboration et al., 2025) has obtained in the radio and sub-mm range.
acknowledgments
The authors acknowledge an anonymous reviewer for their constructive comments. AC acknowledges the UPES Seed grant (UPES/R&D/SoAE/25062025/23) for partially supporting this research. The research of A.C., S.H., and S.S.H. were supported by the Canadian Space Agency (CSA) and the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grants and the Canada Research Chairs programs. The Basic Science Research Program has supported K.H. through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (2016R1A5A1013277 (K.H. and A.C.), 2020R1A2C1007219, and RS-2025-23323627 (K.H.)). K.H. thanks Andrew Mummery for fruitful discussions. AJ acknowledges support from the Fondecyt fellowship (Proyecto 3230303). NK acknowledge funds from European Union - Next Generation EU, Mission 4 Component 1 CUP C53D23001330006 and INAF Large Grant 2023 BLOSSOM F.O. 1.05.23.01.13. Research work at the Physical Research Laboratory, Ahmedabad, is funded by the Department of Space, Government of India. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France.